# IPR-1: Interactive Physical Reasoner

### v1

@@ 5.2 Why is PhysCode Necessary
5.3 Playing in Diverse Physical Worlds of Games
@@
We assess every model on the three hierarchical objectives, instantiating level-specific training or prompting. Further implementation details are provided in the supplementary. The key results are reported in Tab. 2. Takeaways are below the table.
5.4 Zero-shot Transferring to Unseen Games
@@ 6 Discussion
We study an interactive physical reasoning paradigm in which a general-purpose VLM reasons in language, acts through a physics-centric latent interface (PhysCode), and is reinforced by imagined rewards from a world model, asking whether such agents can internalize physical and causal regularities from heterogeneous games and show clear scaling as experience grows. From this perspective, latent-action world models (e.g. Genie, UniVLA [bruce2024genie, bu2025univlalearningacttaskcentric]) learn discrete action abstractions and latent dynamics for controllable rollouts; imagination-based control methods (e.g. Dreamer, V-JEPA2-AC [hafner2024masteringdiversedomainsworld, assran2025vjepa2selfsupervisedvideo]) optimize policies inside learned world models over device-level actions; and large-scale VLM-based game agents (e.g. Game-TARS [wang2025gametarspretrainedfoundationmodels]) scale vision–language–action models with massive human demonstrations and auxiliary multimodal tasks. Yet, from a physics-centric perspective, these approaches do not explicitly organize actions by shared physical mechanisms across hundreds of games or align VLM’s reasoning ability with prediction competence in a common latent space. IPR combines their advantages to study how physical knowledge and transfer emerge under the unified Survival-Curiosity-Utility evaluation, though it is still limited to game environments and short-horizon imagination, leaving real-world transfer and longer-horizon reasoning to future work.

### v2

@@ 5.2 Why is PhysCode Necessary
Table 2: Comprehensive comparison across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. “@” denotes the optimization objective. Scores are normalized individually for each game, scaled between random (0) and human (11) benchmarks. Mean is the average of these normalized scores, indicating overall competence. Avg. Rank is the average relative rank among 30 methods across all games (lower is better). Ratio@Top-3(%) is the proportion of games where the method ranks within the top-3. Our IPR demonstrates robust performance across all metrics.
Methods [Uncaptioned image]Survival [Uncaptioned image]Curiosity [Uncaptioned image]Utility Overall
(Overall) (Robustness) (Competitiveness) (Overall) (Robustness) (Competitiveness) (Overall) (Robustness) (Competitiveness) Avg.
Mean $\uparrow$ Avg. Rank $\downarrow$ Ratio@Top-3(%) $\uparrow$ Mean $\uparrow$ Avg. Rank $\downarrow$ Ratio@Top-3(%) $\uparrow$ Mean $\uparrow$ Avg. Rank $\downarrow$ Ratio@Top-3(%) $\uparrow$ Rank $\downarrow$
Control Group
Random 0.000 $16.2$ $6.7$ 0.000 $18.1$ $3.0$ 0.000 $12.3$ $12.8$ $26.9$
Human 1.000 $5.7$ $46.3$ 1.000 $7.9$ $14.0$ 1.000 $2.9$ $61.6$ $2.8$
Imitation Learning (IL) Group
ACT-BC 0.088 $14.3$ $17.1$ 0.793 $15.1$ $12.8$ 0.255 $12.0$ $13.4$ $16.6$
Qwen3-VL-8B-BC 0.099 $12.9$ $14.0$ 0.812 $12.8$ $9.1$ 0.368 $9.6$ $12.8$ $13.3$
Reinforcement Learning (RL) Group
PPO@survival 0.125 $14.0$ $14.0$ 0.233 $16.5$ $3.7$ 0.588 7.3 30.5 $12.0$
PPO@curiosity 0.114 $14.9$ $11.6$ 0.190 $17.3$ $2.4$ 0.609 6.9 29.3 $14.8$
PPO@utility 0.120 $15.0$ $12.2$ 0.220 $16.8$ $3.0$ 0.534 $8.0$ $25.6$ $14.7$
DQN@survival 0.121 $14.4$ $15.9$ 0.856 $14.4$ $8.5$ 0.497 $10.8$ $15.2$ $12.2$
DQN@curiosity 0.131 $13.2$ 18.3 0.772 $13.4$ $7.9$ 0.424 $10.9$ $15.9$ $10.6$
DQN@utility 0.125 $13.7$ $16.5$ 0.620 $14.2$ $4.9$ 0.445 $10.8$ $17.1$ $11.4$
World Model Group
DreamerV3@survival 0.102 $15.8$ $15.2$ 1.120 12.5 $16.5$ 0.298 $11.3$ $16.5$ $13.1$
DreamerV3@curiosity 0.108 $14.5$ $17.7$ 1.161 $13.1$ $14.0$ 0.235 $10.0$ $20.1$ $10.7$
DreamerV3@utility 0.097 $14.9$ $17.7$ 0.964 $15.4$ $11.0$ 0.139 $11.4$ $18.3$ $15.4$
V-JEPA2@survival 0.102 $17.4$ $4.9$ 1.150 $15.6$ 17.7 0.191 $13.9$ $16.5$ $18.3$
V-JEPA2@curiosity 0.100 $17.8$ $2.4$ 1.402 $15.6$ $16.5$ 0.146 $14.0$ $11.6$ $20.8$
V-JEPA2@utility 0.102 $17.5$ $1.8$ 1.136 $14.5$ 22.6 0.152 $14.1$ $11.6$ $20.2$
GenieRedux@survival 0.108 $13.7$ $15.9$ 1.198 12.5 $11.0$ 0.128 $12.7$ $12.8$ $14.2$
GenieRedux@curiosity 0.104 $14.3$ $14.0$ 1.374 $12.5$ $9.8$ 0.100 $12.8$ $12.8$ $16.1$
GenieRedux@utility 0.110 $13.7$ $16.5$ 1.248 12.4 $14.6$ 0.122 $13.5$ $14.6$ $12.4$
Multimodal Large Language Model (MLLM) Group
GPT-4o@survival 0.108 12.6 $13.4$ 0.039 $17.2$ $0.6$ 0.302 $9.2$ $19.5$ $16.4$
GPT-4o@curiosity 0.079 $16.8$ $11.6$ 0.368 $15.3$ $5.5$ 0.186 $10.6$ $17.7$ $19.4$
GPT-4o@utility 0.087 $15.8$ $10.4$ 0.319 $14.6$ $3.7$ 0.337 $10.0$ $17.1$ $18.8$
GPT-5@survival 0.140 10.5 24.4 0.127 $18.3$ $1.8$ 0.263 $8.0$ $23.8$ $13.3$
GPT-5@curiosity 0.093 $15.3$ $12.2$ 0.298 $16.4$ $7.3$ 0.333 $9.8$ $16.5$ $17.9$
GPT-5@utility 0.108 $15.2$ $11.0$ 0.185 $16.5$ $0.6$ 0.371 7.8 26.2 $16.8$
Qwen3-VL-30B-A3B@survival 0.091 $14.3$ $11.0$ 0.325 $23.0$ $0.0$ 0.289 $12.0$ $14.0$ $22.7$
Qwen3-VL-30B-A3B@curiosity 0.086 $15.8$ $11.6$ 0.878 $20.5$ $2.4$ 0.155 $11.7$ $15.2$ $22.4$
Qwen3-VL-30B-A3B@utility 0.108 $13.5$ $12.2$ 0.528 $21.3$ $4.9$ 0.285 $11.6$ $14.6$ $17.6$
Interactive Physical Reasoner
Qwen3-VL-8B w/o IPR 0.105 $13.7$ $14.0$ 0.325 $15.0$ $4.3$ 0.176 $11.6$ $12.8$ $18.2$
Qwen3-VL-8B w/ IPR 0.252 2.6 72.0 1.173 $13.1$ $13.4$ 0.493 $8.5$ $22.0$ 4.9
(IPR ranking w/o control group) (1/28) (1/28) (1/28) (5/28) (6/28) (7/28) (5/28) (6/28) (6/28) (4.9/28)
Key Takeaways across [Uncaptioned image]Survival, [Uncaptioned image]Curiosity, and [Uncaptioned image]Utility Prediction-based Methods (WM). Strong at [Uncaptioned image], but weaker at [Uncaptioned image] and [Uncaptioned image]. Trained on broad exploratory trajectories, latent rollouts broaden coverage and reveal dynamics, but tend to imitate visually-alike futures rather than reliably pursue goals. So prediction is useful as a look-ahead prior for risk and candidate actions. RL-based Methods (PPO, DQN). Strong at [Uncaptioned image]and [Uncaptioned image] when rewards are well-shaped, but weaker on [Uncaptioned image] and tasks without explicit goals. Reward gradients enable effective credit assignment under the right signal, yet sparsity and partial observability induce instability and interface overfitting, so RL works best as an optimization method. Experience-based Methods (Behavior Cloning). Strong at human-like [Uncaptioned image], but weaker on [Uncaptioned image] and [Uncaptioned image]. Deliberately imitate human trajectories and thus excel at low-risk survival, but struggle once tasks require precise control or exploration, and their performance depends strongly on the coverage and quality of the demonstrations. Reasoning-based Pretrained VLMs. Strong at goal-conditioned [Uncaptioned image]and [Uncaptioned image]; weaker on [Uncaptioned image]. They excel at instruction-driven reasoning but cannot predict consequences in the visual state space, so they work best as high-level reasoners that need auxiliary prediction modules for outcome-aware decisions. Interactive Physical Reasoner (Ours). Robust across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. We combine the strengths of all three paradigms: VLMs provide goal-driven causal reasoning, the world model supplies rollout prediction, and RL optimizes decisions using imagined rewards, yielding consistently strong performance across all three levels. Summary. Prediction-based world models understand dynamics but cannot reliably plan toward long-horizon goals, while reasoning-based VLMs can plan semantically but lack grounded prediction of physical outcomes. IPR combines them by using WM rollouts as physical priors and VLM reasoning to select and pursue feasible futures, surpassing GPT-5 with an 8B backbone.
5.3 Playing in Diverse Physical Worlds
@@
We assess every model on the three hierarchical objectives, instantiating level-specific training or prompting. Further implementation details are provided in the supplementary. The key results are reported in Tab. 2. Takeaways are summarized below the table.
5.4 Zero-shot Transfer to Unseen Games
@@ 6 Discussion
We study an interactive physical reasoner paradigm in which a general-purpose VLM reasons in language, acts through a physics-centric latent interface (PhysCode), and is reinforced by imagined rewards from a world model, asking whether such agents can internalize physical and causal regularities from heterogeneous games and show clear scaling as experience grows. From this perspective, latent-action world models (e.g. Genie, UniVLA [bruce2024genie, bu2025univlalearningacttaskcentric]) learn discrete action abstractions and latent dynamics for controllable rollouts; imagination-based control methods (e.g. Dreamer, V-JEPA2-AC [hafner2024masteringdiversedomainsworld, assran2025vjepa2selfsupervisedvideo]) optimize policies inside learned world models over device-level actions; and large-scale VLM-based game agents (e.g. Game-TARS [wang2025gametarspretrainedfoundationmodels]) scale vision–language–action models with massive human demonstrations and auxiliary multimodal tasks. Yet, from a physics-centric perspective, these approaches do not explicitly organize actions by shared physical mechanisms across hundreds of games or align VLM’s reasoning ability with prediction competence in a common latent space. IPR combines their advantages to study how physical knowledge and transfer emerge under the unified Survival-Curiosity-Utility evaluation, though it is still limited to game environments and short-horizon imagination, leaving real-world transfer and longer-horizon reasoning to future work.